• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 06:22:58 on July 27, 2006

    Entry number 175534

    keyword=Shift Summary (owl...up to 6am)

    12:40 Beam in Hall. They can get current up to 17 uA without any problems. Did some harp scans, and beam size was good - roughly 300 x 300 microns (see entry 175492). Set raster size again (see other halog entries about that).

    We have a quad in each spectrometer that needs resetting, but we are unable to contact a tech to go reset ... not a big disaster since we can commission the radiator; we get more useful information if the HRS-L was up and running (since we could also acquire our first gamma-d spectra and check our background subtraction procedure), however the radiator commissioning can proceed without the spectrometers - so we'll have the quads reset first thing in the morning.

    Ron Gilman comes in for radiator commissioning. After acquiring a lot of confusing data, he went down to the hall and discovered the controller was not placing the foils in the corect positions (putting halfway between...ie. in empty spaces!). Reboot the IOC that controls the radiator motion, and re-calibrate positions.

    Now radiator seems to move as it should (see Gilman's halog). Radiator commissioning progresses now smoothly.

    Nothing else can be done until quads are reset. Stay with beam on radiator, LD2 target in.